# A logic of co-valuations

### v1

@@ 4.3. Compact structures
The topology $\tau_{V}$ is generated by its basis $\bigcup_{v\in V}[v]$. If $(M,V)$ is a compact $\omega$-structure, then this is a cap-basis.
If $\tau_{V}$ is metrizable (connected, 0-dimensional, etc.), we call $(M,V)$ a metrizable (connected, 0-dimensional, etc.) structure. For compact structures $(M,V)$, $(N,W)$ in the same signature, a bijection $f:M\to N$ is called a topological isomorphism if it preserves relations, and is a homeomorphism of $(M,\tau_{V})$ and $(N,\tau_{W})$.
@@
Let $(M,\tau)$ be a compact $T_{1}$ space, let $\{R^{M}_{\alpha}\}$ be a family of closed relations on $M$, and let $\mathcal{C}$ be a family of minimal open covers of $M$ such that for any $A,B\in\mathcal{C}$ there is $C\in\mathcal{C}$ that fragments $A$, $B$, and every open cover is refined by some $C\in\mathcal{C}$. Let $V$ be the family generated by of all co-valuations defined as in Remark 4.2, using covers in $\mathcal{C}$. Then $(M,\{R^{M}_{\alpha}\},V)$ is a compact structure in signature $L=\{R_{\alpha}\}$ such that $\tau_{V}=\tau$.
@@
We only need to verify Axiom (II). Fix $v\in V$, a relation $R^{M}_{\alpha}\subseteq M^{m}$, an $m$-tuple $\bar{x}=(x_{1},\ldots,x_{m})$ in $[v]$, and an $m$-tuple $\bar{a}$ in $M$ such that $a_{k}\in x_{k}$, and $R^{M}_{\alpha}(\bar{a})$. Because each $x_{k}$ is open, we can find open $z^{\prime}_{k}$ such that $a_{k}\in z^{\prime}_{k}\subseteq x_{k}$ and $a_{l}\not\in z^{\prime}_{k}$ if $a_{l}\neq a_{k}$. Fix a minimal open cover $Z^{\prime}$ of $M\setminus\bigcup_{k}z_{k}$ consisting of sets $z^{\prime}$ such that $z^{\prime}\cap\{a_{k}\}=\emptyset$, and $Z^{\prime\prime}=Z^{\prime}\cup\{z^{\prime}_{k}\}$ refines $v$. Finally, fix $w^{\prime}\in V$ that refines $Z^{\prime\prime}$, and fragments $v$. One can verify that $w=C_{\geq^{Z^{\prime\prime}}_{w^{\prime}}}(w^{\prime})$, $\bar{a}$ are witnesses of Axiom (II) for $v$, $\bar{x}$, $R^{M}_{\alpha}$. ∎
@@
Let $(M,\tau)$ be a second-countable compact $T_{1}$ space, let $\{R^{M}_{\alpha}\}$ be a family of closed relations on $M$, and let $\mathbb{P}_{i}$, $i\in\mathbb{N}$, be a sequence of finite minimal open covers of $M$ such that every $\mathbb{P}_{i}$ consolidates $\mathbb{P}_{i+1}$, and every open cover of $M$ is refined by some $\mathbb{P}_{i}$. Then $(M,\{R^{M}_{\alpha}\},V)$, where $V=\langle\{v_{i}\}\rangle$, and $v_{i}$ are defined as in Remark 4.2, is a compact $\omega$-structure in signature $L=\{R_{\alpha}\}$ such that $\tau_{V}=\tau$.
@@
Let $(M,V)$ be a structure in signature $L$, and let $(v_{i})$ be an $\omega$-chain of admissible co-valuations. Let $\mathbb{P}=\bigsqcup[v_{i}]$ be ordered by inclusion. We define a compact $\omega$-stucture $(N,W)$, called the compact $\omega$-structure induced by $(v_{i})$, as follows. Let $N={\mathsf{S}}\mathbb{P}$. Define co-valuations $w_{i}\in{\rm{CoV}}(X_{v_{i}},N)$ induced by $(v_{i})$, by
@@
and put $W=\langle\{w_{i}\}\rangle$. Finally, for $R\in L$ of arity $m$, $S_{k}\in N$, $k\leq m$, define
@@
for all $i\in\mathbb{N}$, and $x_{k}$, $k\leq m$, such that $x_{k}\in S_{k}$. Clearly, all $R^{N}$ are closed in $\tau_{W}$. By Corollary 4.14, $(N,W)$ is a compact $\omega$-structure.
@@ 4.4. Properties of compact structures
Let $(M,V)$ be a compact structure, let $X$ be a context, let $x_{0}\in X$, $X_{0}\subseteq X$, and $v\in V^{X}$. The following properties of $v$ can be expressed by a single formula:
@@
To see the other direction, suppose that $c=\bigcap[v]_{X_{0}}\neq\emptyset$. Fix a minimal open cover $C$ of $M$ extending $\{c\}$, and let $w^{\prime}\in V^{Z}$ be a fragmentation of $v$ such that $w^{\prime}]$ refines $C$. Fix $y_{0}\in Z$ such that $y_{0}]_{w^{\prime}}\subseteq c$. Note that, by minimality of $[v]$, $c\neq[x]$, for $x\in X$. Finally, let $w\in V^{Y}$ be a consolidation of $w^{\prime}$ determined by the context $Y=X\cup\{y_{0}\}$. Then $w$ witnesses that $M\models_{v}\exists_{\geq}\top$.
@@
To see the other direction, fix distinct $a_{0},a_{1}\in[x_{0}]_{v}$. Fix open $c_{0},c_{1}\subseteq[x_{0}]_{v}$ such that $a_{0}\in c_{0}\setminus c_{1}$, $a_{1}\in c_{1}\setminus c_{0}$, a minimal open cover $C$ of $M$ extending $\{c_{0},c_{1}\}$ such that $a_{0},a_{1}\not\in\bigcup C\setminus\{c_{0},c_{1}\}$, and let $w^{\prime}\in V^{Z}$ be a fragmentation of $v$ such that $w^{\prime}]$ refines $C$. Fix $y_{0},y_{1}\in Y$ such that $a_{0}\in[y_{0}]_{w^{\prime}}$, $a_{1}\in[y_{1}]_{w^{\prime}}$; in particular $y_{0}\neq y_{1}$. Finally, let $w$ be a consolidation of $w^{\prime}$ determined by $Y$. Then $w$ witnesses that $M\models_{v}\exists_{\geq}\top$. ∎
@@
Let $X$ be a context, and let $G\subseteq X\times X$ be a reflexive graph. By Proposition 4.17(1), we can express by a single formula in the context $X$ the fact that $G(v)=G$, for every $(M,V)$, and $v\in V^{X}$. We denote this formula by $\mbox{OverlapGraph}_{G}(X)$.
@@
Suppose that a compact $\omega$-structure $(M,V)$ omits all the types $A_{n}(X_{n})$. As we observed above, if for some $v\in V^{X}$, $G\subseteq X\times X$, and $\geq\in\mbox{Star}_{G}(X,Y)$, $w\in V^{Y}$ witnesses that
@@
Later we will see that more can be proved about certain metrizable $\omega$-structures topologically isomorphic with the arc, the ordered arc, and with the pseudo-arc.
@@ 5. Elements of model theory
Let $D$ be an ultrafiler on a set $I$, and let $(M_{i},V_{i})$, $i\in I$, be proto-structures in signature $L$. For a context $X$, the ultraproduct $\prod_{D}v_{i}\in{\rm{CoV}}(X,\prod_{D}M_{i})$ of co-valuations $v_{i}\in V_{i}^{X}$ on the ultraproduct $\prod_{D}M_{i}$ in $X$ is defined by
$(x,(a_{i}))\in\prod_{D}v_{i}\iff(x,a_{i})\in v_{i}\text{ for $D$-many $i$}.$
We denote the family of all ultraproducts of admissible co-valuations by $\prod_{D}V_{i}$. It is immediate to verify that
$(a_{i})\in[x]\pi_{\prod_{D}v_{i}}\iff a_{i}\in[x]\pi_{v_{i}}\text{ for $D$-many $i$},$
@@
In particular, $(\prod_{D}M_{i},\prod_{D}V_{i})$ is a proto-structure. By Proposition 4.7, if for proto-structures $M_{i}$, admissible co-valuations satisfy Axiom (I), then admissible co-valuations satisfy Axiom (I) for $(\prod_{D}M_{i},\prod_{D}V_{i})$. By the above observations, if $w_{i}$, $\bar{a}_{i}$ are witnesses for $R$, $v_{i}$, $\bar{x}$, then $\prod_{D}w_{i}$, $\prod_{D}\bar{a}_{i}$ are witnesses for $R$, $\prod_{D}v_{i}$, $\bar{x}$. Hence, $(\prod_{D}M_{i},\prod_{D}V_{i})$ is a structure, i.e., the class of structures is closed under ultraproducts, and the following holds:
@@
Let $D$ be an ultrafilter on a set $I$, let $(M_{i},V_{i})$, $i\in I$, be a family of structures in signature $L$, and let $v_{i}\in V_{i}$ be co-valuations in context $X$, $i\in I$. For every formula $\phi(X)$,
@@
The proof is by induction on the complexity of formulas. The only non-trivial case is $\exists^{\scriptscriptstyle{f}}_{\geq}\phi$. To see the direction from left to right, take some admissible co-valuation, which is of the form $\prod_{D}w_{i}$, such that $M\models_{\prod_{D}w_{i}}\phi$, and $\prod_{D}v_{i}$ consolidates $\prod_{D}w_{i}$ with pattern $\geq$. Since $v_{i}$ consolidates $w_{i}$ with pattwern $\geq$, for $D$-many $i$, by the induction hypothesis, for $D$-many $i$, we have $M_{i}\models_{w_{i}}\phi$. Then we have $M_{i}\models_{v_{i}}\exists^{\scriptscriptstyle{f}}_{\geq}\phi$ for $D$-many ii. To see the other direction, suppose $M_{i}\models_{v_{i}}\exists^{\scriptscriptstyle{f}}_{\geq}\phi$ for $D$-many ii. For those $i$, take $w_{i}$ such that $v_{i}$ consolidates $w_{i}$ with pattern $\geq$ and $M_{i}\models_{w_{i}}\phi$, and let $w_{i}$ be an arbitrary admissible co-valuation of $M_{i}$ in $X$ otherwise. Then $\prod_{D}v_{i}$ consolidates $\prod_{D}w_{i}$ with pattern $\geq$, and $\prod_{D}M_{i}\models_{\prod_{D}w_{i}}\phi$ by the induction hypothesis, so we obtain $\prod_{D}M_{i}\models_{\prod_{D}w_{i}}\exists^{\scriptscriptstyle{f}}_{\geq}\phi$. ∎
@@
A type is consistent iff its every finite subset is consistent.
@@ 5.2. The downward Löwenheim-Skolem theorem
Fix $v_{0}\in V$. Using Axioms (I) and (II), construct an $\omega$-chain $(v_{i})$ in $V$ such that, for $V^{\prime}=\langle\{v_{i}\}\rangle$,
@@
Let $\mathbb{P}=\bigsqcup[v_{i}]$ be ordered by inclusion, and let $({\mathsf{S}}\mathbb{P},W)$ be the compact $\omega$-structure induced by $(v_{i})$, where $W=\langle\{w_{i}\}\rangle$, and $(w_{i})$ is the core induced by $(v_{i})$. We claim that, for every $i\in\mathbb{N}$,
@@
Fix $R\in L$ be of arity $m$. Suppose that $N\models_{w_{i}}R(\bar{x})$, for $\bar{x}=(x_{k})$. This means that there are $S_{k}\in(x_{k})^{\in}$, $k\leq m$, such that $R^{N}((S_{k}))$, so $M\models_{v_{i}}R(\bar{x})$. On the other hand, if $M\models_{v_{i}}R(\bar{x})$, using Axiom (II), we fix witnesses $v\in(V^{\prime})^{Y}$, $\bar{y}\in Y^{m}$, $\bar{a}$ for $v_{i}$, $\bar{x}$, $R$, i.e., $\bar{a}\in[\bar{y}]_{\pi_{v}}$, $R^{M}(\bar{a})$, and $\bar{x}\geq^{v_{i}}_{v}\bar{y}$. For $k\leq m$, consider $S^{\prime}_{k}=\{x\in\mathbb{P}:a_{k}\in x\}$. As every $v_{i^{\prime}}]$, $i^{\prime}\in\mathbb{N}$, is a cover of $M$, every level $\mathbb{P}_{i^{\prime}}$ overlaps $S^{\prime}_{k}$. Hence, by Remark 3.7, $S^{\prime}_{k}$ is a selector, and, by Lemma 3.8, there is a minimal selector $S_{k}\subseteq S^{\prime}_{k}$. Let $j\in\mathbb{N}$ be such that $v_{j}$ fragments $v$. Observe that for every $y\in[v_{j}]$ we have that $y\in S^{\prime}_{k}$ implies $x_{k}\geq^{v_{i}}_{v}\circ\geq^{v_{i}}_{v_{j}}y$; in particular, it must hold that $S_{k}\in(x_{k})^{\in}$. Finally, $\bar{a}$ ensures that for every $i^{\prime}\in\mathbb{N}$, if $x^{\prime}_{k}\in S_{k}$, $k\leq m$, then $M\models_{v_{i^{\prime}}}R((x^{\prime}_{k}))$. Thus, $R^{N}((S_{k}))$, and $N\vDash_{w_{i}}R(\bar{x})$.
To finish the proof, we proceed by induction on the complexity of formulas. Clearly, it suffices to consider the case $\exists^{\scriptscriptstyle{f}}_{\geq}\phi$. Let $i\in\mathbb{N}$, and $\phi$ be such that $M\models_{v_{i}}\exists^{\scriptscriptstyle{f}}_{\geq}\phi$, and let $v\in V$ witness it, i.e., $v_{i}=C_{\geq}(v)$ and $M\models_{v}\phi$. By (3), we can assume that $v\in V^{\prime}$, i.e., there are $j\geq i$, $\geq_{1}$ and $\geq_{2}$ such that $v_{i}=C_{\geq_{1}}(v_{j})$, $v=C_{\geq_{2}}(v_{j})$. By the inductive assumption, $N\models_{C_{\geq_{2}}(w_{j})}\phi$. By Corollary 3.16, $C_{\geq_{1}}(w_{j})=w_{i}$, and $C_{\geq_{1}}(w_{j})$ consolidates $C_{\geq_{2}}(w_{j})$ with pattern $\geq$, i.e., $N\models_{w_{i}}\exists^{\scriptscriptstyle{f}}_{\geq}\phi$. By Corollary 3.16 again, the other implication can be proved exactly in the same way.

### v2

@@ 4.3. Compact structures
By Proposition 3.7, the topology $\tau_{V}$ is generated by its basis $\bigcup_{v\in V}[v]$. In the case that $(M,V)$ is a compact $\omega$-structure, by Remark 3.11, this is a cap-basis.
If $\tau_{V}$ is metrizable (connected, 0-dimensional, etc.), we call $(M,V)$ a metrizable (connected, 0-dimensional, etc.) structure. For compact structures $(M,V)$, $(N,W)$ in the same signature, a bijection $f:M\to N$ is a topological isomorphism if it preserves relations, and is a homeomorphism of $(M,\tau_{V})$ and $(N,\tau_{W})$.
@@
Let $(M,\tau)$ be a compact $T_{1}$ space, let $\{R^{M}_{\alpha}\}$ be a family of closed relations on $M$, and let $\mathcal{C}$ be a family of minimal open covers of $M$ such that for any $A,B\in\mathcal{C}$ there is $C\in\mathcal{C}$ that fragments $A$, $B$, and every open cover of $M$ is refined by some $C\in\mathcal{C}$. Let $V$ be the family generated by of all co-valuations defined as in Remark 4.2, using covers in $\mathcal{C}$. Then $(M,\{R^{M}_{\alpha}\},V)$ is a compact structure in signature $L=\{R_{\alpha}\}$ such that $\tau_{V}=\tau$.
@@
We only need to verify Axiom (II). Fix $v\in V$, a relation $R^{M}_{\alpha}\subseteq M^{m}$, an $m$-tuple $\bar{x}=(x_{1},\ldots,x_{m})$ in $[v]$, and an $m$-tuple $\bar{a}=(a_{1},\ldots,a_{m})$ in $M$ such that $a_{k}\in x_{k}$ and $R^{M}_{\alpha}(\bar{a})$. Because each $x_{k}$ is open, we can find open $z^{\prime}_{k}$ such that $a_{k}\in z^{\prime}_{k}\subseteq x_{k}$ and $a_{l}\not\in z^{\prime}_{k}$, if $a_{l}\neq a_{k}$. Fix a minimal open cover $Z^{\prime}$ of the compact subspace $M\setminus\bigcup_{k}z_{k}$ consisting of sets $z^{\prime}$ such that $z^{\prime}\cap\{a_{k}\}=\emptyset$, $k\leq m$, and $Z^{\prime\prime}=Z^{\prime}\cup\{z^{\prime}_{k}\}$ refines $v$. Finally, fix $w^{\prime}\in V$ that refines $Z^{\prime\prime}$, and fragments $v$. One can easily verify that $w=C_{\geq^{Z^{\prime\prime}}_{w^{\prime}}}(w^{\prime})$, $\bar{a}$ are witnesses of Axiom (II) for $v$, $\bar{x}$, $R^{M}_{\alpha}$. ∎
@@
Let $(M,\tau)$ be a second-countable compact $T_{1}$ space, let $\{R^{M}_{\alpha}\}$ be a family of closed relations on $M$, and let $\mathbb{P}_{i}$, $i\in\mathbb{N}$, be a sequence of minimal open covers of $M$ such that every $\mathbb{P}_{i}$ consolidates $\mathbb{P}_{i+1}$, and every open cover of $M$ is refined by some $\mathbb{P}_{i}$. Then $(M,\{R^{M}_{\alpha}\},V)$, where $V=\langle\{v_{i}\}\rangle$, and $v_{i}$ are defined as in Remark 4.2, is a compact $\omega$-structure in signature $L=\{R_{\alpha}\}$ such that $\tau_{V}=\tau$.
@@
Let $(M,V)$ be a structure in signature $L$, and let $(v_{i})$ be an $\omega$-chain of admissible co-valuations. Let $\mathbb{P}$ be the $\omega$-chain poset $(\bigsqcup[v_{i}],\supseteq)$. We define a compact $\omega$-stucture $(N,W)$, called the compact $\omega$-structure induced by $(v_{i})$, as follows. We put $N={\mathsf{S}}\mathbb{P}$, define co-valuations $w_{i}\in{\rm{CoV}}(X_{v_{i}},N)$ induced by $v_{i}$, by
@@
(i.e., $w_{i}$ is defined essentially as in Remark 4.2, using $v_{i}]^{\in}$), and put $W=\langle\{w_{i}\}\rangle$. Finally, for $R\in L$ of arity $m$, $S_{k}\in N$, $k\leq m$, we define
@@
for all $i\in\mathbb{N}$, and $x_{k}$, $k\leq m$, such that $x_{k}\in S_{k}$. Clearly, all $R^{N}$ are closed in $\tau_{W}$. By Proposition 3.13, and Corollary 4.16, $(N,W)$ is a compact $\omega$-structure.
@@ 4.4. Properties of compact structures
Let $(M,V)$ be a compact structure, let $X$ be a context, let $x_{0}\in X$, $X_{0}\subseteq X$, where $|X_{0}|>1$, and $v\in V^{X}$. The following properties of $v$ can be expressed by a single formula:
@@
To see the other direction, suppose that $c=\bigcap[v]_{X_{0}}\neq\emptyset$. Fix a minimal open cover $C$ of $M$ extending $\{c\}$, and let $w^{\prime}\in V^{Z}$ be a fragmentation of $v$ such that $w^{\prime}]$ refines $C$. Fix $y_{0}\in Z$ such that $y_{0}]_{w^{\prime}}\subseteq c$. Note that, by minimality of $[v]$, for every $x\in X$, there is $z\in Z$ such that $z\neq y_{0}$, $x]_{v}\supseteq[z]_{w^{\prime}}$, and $x^{\prime}]_{v}\supseteq[z]_{w^{\prime}}$ implies $x=x^{\prime}$. Finally, let $\succeq\in{\rm{CoV}}(Y,Z)$ be a pattern defined by $y\succeq z$ iff $y=z=y_{0}$ or ($z\neq y_{0}$ and $x]_{v}\supseteq[z]_{w^{\prime}}$). Then $w=C_{\succeq}(w^{\prime})$ witnesses that $M\models_{v}\exists_{\geq}\top$.
@@
To see the other direction, fix distinct $a_{0},a_{1}\in[x_{0}]_{v}$. Fix open $c_{0},c_{1}\subseteq[x_{0}]_{v}$ such that $a_{0}\in c_{0}\setminus c_{1}$, $a_{1}\in c_{1}\setminus c_{0}$, a minimal open cover $C$ of $M$ extending $\{c_{0},c_{1}\}$ such that $a_{0},a_{1}\not\in\bigcup C\setminus\{c_{0},c_{1}\}$, and let $w^{\prime}\in V^{Z}$ be a fragmentation of $v$ such that $w^{\prime}]$ refines $C$. Fix $y_{0},y_{1}\in Y$ such that $a_{0}\in[y_{0}]_{w^{\prime}}$, $a_{1}\in[y_{1}]_{w^{\prime}}$; in particular $y_{0}\neq y_{1}$. Finally, let $w$ be a consolidation of $w^{\prime}$ determined by $Y$ as in (1). Then $w$ witnesses that $M\models_{v}\exists_{\geq}\top$. ∎
@@
Let $X$ be a context, and let $G\subseteq X\times X$ be a reflexive graph. By Proposition 4.19(1), there is a single formula in the context $X$ that expresses the fact that $G(v)=G$, for every $(M,V)$, and $v\in V^{X}$. We denote this formula by $\mbox{OverlapGraph}_{G}(X)$.
@@
Suppose that a compact $\omega$-structure $(M,V)$ omits all the types $A_{n}(X_{n})$. As it was observed above, if for some $v\in V^{X}$, $G\subseteq X\times X$, and $\geq\in\mbox{Star}_{G}(X,Y)$, $w\in V^{Y}$ witnesses that
@@
Later we will see that more can be proved about certain metrizable $\omega$-structures topologically isomorphic with the arc, the ordered arc, or with the pseudo-arc.
@@ 5. Elements of model theory
Let $D$ be an ultrafiler on a set $I$, and let $(M_{i},V_{i})$, $i\in I$, be proto-structures in signature $L$. Let $\prod_{D}M_{i}$ be the ultraproduct of first-order structures $M_{i}$. For a context $X$, the ultraproduct $\prod_{D}v_{i}\in{\rm{CoV}}(X,\prod_{D}M_{i})$ of co-valuations $v_{i}\in V_{i}^{X}$ on $\prod_{D}M_{i}$ in $X$ is defined by
$(x,[(a_{i})])\in\prod_{D}v_{i}\iff(x,a_{i})\in v_{i}\text{ for $D$-many $i$}.$
Observe that this definition does not depend on the choice of the representative $(a_{i})\in\prod M_{i}$. We denote the family of all ultraproducts of admissible co-valuations by $\prod_{D}V_{i}$. It is immediate to verify that
$(a_{i})]\in[x]_{\pi_{\prod_{D}v_{i}}}\iff a_{i}\in[x]_{\pi_{v_{i}}}\text{ for $D$-many $i$},$
@@
In particular, $(\prod_{D}M_{i},\prod_{D}V_{i})$ is a proto-structure. By Proposition 4.9, if for proto-structures $M_{i}$, admissible co-valuations satisfy Axiom (I), then admissible co-valuations satisfy Axiom (I) for $(\prod_{D}M_{i},\prod_{D}V_{i})$. By the above observations, if $w_{i}$, $\bar{a}_{i}$ are witnesses for $R$, $v_{i}$, $\bar{x}$, then $\prod_{D}w_{i}$, $\prod_{D}\bar{a}_{i}$ are witnesses for $R$, $\prod_{D}v_{i}$, $\bar{x}$. Hence, $(\prod_{D}M_{i},\prod_{D}V_{i})$ is a structure, if all $(M_{i},V_{i})$ are structures. Thus, we have:
Proposition 5.1.
The class of proto-structures, and the class of structures are closed under ultraproducts.
@@
Let $D$ be an ultrafilter on a set $I$, let $(M_{i},V_{i})$, $i\in I$, be a family of proto-structures in signature $L$, and let $v_{i}\in V_{i}$ be co-valuations in context $X$, $i\in I$. For every formula $\phi(X)$,
@@
The proof is by induction on the complexity of formulas. First, observe that, for every $R\in L$, $\prod_{D}M\models_{\prod_{D}v_{i}}R(\bar{x})$ iff there exist $D$-many $i$, and $\bar{a}_{i}\in[\bar{x}]_{v_{i}}$ such $R^{M_{i}}(\bar{a}_{i})$. The only non-trivial inductive step is $\exists^{\scriptscriptstyle{f}}_{\geq}\phi$, where $\phi$ is a formula in context $Y$, and $\geq\in{\rm{CoV}}(X,Y)$. To see the direction from left to right, take some admissible co-valuation $\prod_{D}w_{i}\in{\rm{CoV}}(Y,\prod_{D}M_{i})$ such that $\prod_{D}M\models_{\prod_{D}w_{i}}\phi$, and $\prod_{D}v_{i}$ consolidates $\prod_{D}w_{i}$ with pattern $\geq$. Since $v_{i}$ consolidates $w_{i}$ with pattern $\geq$ for $D$-many $i$, by the induction hypothesis, for $D$-many $i$, we have $M_{i}\models_{w_{i}}\phi$. Then we have $M_{i}\models_{v_{i}}\exists^{\scriptscriptstyle{f}}_{\geq}\phi$ for $D$-many ii. To see the other direction, suppose $M_{i}\models_{v_{i}}\exists^{\scriptscriptstyle{f}}_{\geq}\phi$ for $D$-many ii. For those $i$, take $w_{i}$ such that $v_{i}$ consolidates $w_{i}$ with pattern $\geq$ and $M_{i}\models_{w_{i}}\phi$, and let $w_{i}$ be an arbitrary admissible co-valuation of $M_{i}$ in $X$ otherwise. Then $\prod_{D}v_{i}$ consolidates $\prod_{D}w_{i}$ with pattern $\geq$, and $\prod_{D}M_{i}\models_{\prod_{D}w_{i}}\phi$ by the induction hypothesis, so we obtain $\prod_{D}M_{i}\models_{\prod_{D}w_{i}}\exists^{\scriptscriptstyle{f}}_{\geq}\phi$. ∎
@@
A type is consistent iff all its finite subsets are consistent.
@@ 5.2. The downward Löwenheim-Skolem theorem
Fix $v_{0}\in V$. Using Axioms (I) and (II), we construct an $\omega$-chain $(v_{i})$ in $V$ such that, for $V^{\prime}=\langle\{v_{i}\}\rangle$,
@@
Let $\mathbb{P}$ be the $\omega$-chain poset $(\bigsqcup[v_{i}],\supseteq)$. Let $({\mathsf{S}}\mathbb{P},W)$ be the compact $\omega$-structure induced by $(v_{i})$, where $W=\langle\{w_{i}\}\rangle$, and $(w_{i})$ is the core induced by $(v_{i})$. We claim that, for every $i\in\mathbb{N}$,
@@
Fix $R\in L$ be of arity $m$. Suppose that $N\models_{w_{i}}R(\bar{x})$, for $\bar{x}=(x_{k})$. This means that there are $S_{k}\in(x_{k})^{\in}$, $k\leq m$, such that $R^{N}((S_{k}))$, so $M\models_{v_{i}}R(\bar{x})$. On the other hand, if $M\models_{v_{i}}R(\bar{x})$, using Axiom (II), we fix witnesses $v\in(V^{\prime})^{Y}$, $\bar{y}\in Y^{m}$, $\bar{a}$ for $v_{i}$, $\bar{x}$, $R$, i.e., $\bar{a}\in[\bar{y}]_{\pi_{v}}$, $R^{M}(\bar{a})$, and $\bar{x}\geq^{v_{i}}_{v}\bar{y}$. For each $k\leq m$, we consider $S^{\prime}_{k}=\{x\in\mathbb{P}:a_{k}\in x\}$. As every $v_{i^{\prime}}]$, $i^{\prime}\in\mathbb{N}$, is a cover of $M$, every level $\mathbb{P}_{i^{\prime}}$ has non-empty intersection with $S^{\prime}_{k}$. Hence, by Remark 3.8, $S^{\prime}_{k}$ is a selector, and, by Lemma 3.9, there is a minimal selector $S_{k}\subseteq S^{\prime}_{k}$. Let $j\in\mathbb{N}$ be such that $v_{j}$ fragments $v$. Observe that for every $y\in[v_{j}]$ we have that $y\in S^{\prime}_{k}$ implies $x_{k}\geq^{v_{i}}_{v}\circ\geq^{v}_{v_{j}}y$; in particular, it must hold that $S_{k}\in(x_{k})^{\in}$. Finally, $\bar{a}$ ensures that for every $i^{\prime}\in\mathbb{N}$, if $x^{\prime}_{k}\in S_{k}$, $k\leq m$, then $M\models_{v_{i^{\prime}}}R((x^{\prime}_{k}))$. Hence, $R^{N}((S_{k}))$, and $N\vDash_{w_{i}}R(\bar{x})$.
To finish the proof, we proceed by induction on the complexity of formulas. Clearly, the only non-trivial case is $\exists^{\scriptscriptstyle{f}}_{\geq}\phi$. Let $i\in\mathbb{N}$, and $\phi$ be such that $M\models_{v_{i}}\exists^{\scriptscriptstyle{f}}_{\geq}\phi$, and let $v\in V$ witness it, i.e., $v_{i}=C_{\geq}(v)$ and $M\models_{v}\phi$. By (3), we can assume that $v\in V^{\prime}$, i.e., there are $j\geq i$, $\geq_{1}$ and $\geq_{2}$ such that $v_{i}=C_{\geq_{1}}(v_{j})$, $v=C_{\geq_{2}}(v_{j})$. By the inductive assumption, $N\models_{C_{\geq_{2}}(w_{j})}\phi$. By Corollary 3.17, $C_{\geq_{1}}(w_{j})=w_{i}$, and $C_{\geq_{1}}(w_{j})$ consolidates $C_{\geq_{2}}(w_{j})$ with pattern $\geq$, i.e., $N\models_{w_{i}}\exists^{\scriptscriptstyle{f}}_{\geq}\phi$. Using Corollary 3.17 again, the other implication can be proved exactly in the same way.